OpenStudy (anonymous):

Could you please explain: . Find the rate of change of the area of a circle with respect to the radius r. What is the rate when r = 2?

OpenStudy (cwrw238):

rate of change of area = d A / dr = d(pi r^2)/ dr = 2 PI r when r = 2 dA /dr = 4 pi

OpenStudy (cwrw238):

ok? - i've assumed you know calculus

OpenStudy (anonymous):

i have just started learning calc this year but i dont understand other problems like this, i just know the circle ones hahaha

OpenStudy (anonymous):

for example . Find the rate of change of the area of a square with respect to the length z of a diagonal. What is the rate when z = 4?.